OpenStudy (anonymous):

How can I graph the linear equation -x + 8y =15

OpenStudy (the_fizicx99):

Add x to both sides and divide by 8, -x + 8y = 15 +x +x 8y = 15 + x /8 /8 /8 y = 1.875 x/8 x/8 is the same thing as 1/8. which is the slope. The slope is rise / run, so you would rise 1 from 1.875 ( which is the y intercept) and run the the right 8 spots.

OpenStudy (anonymous):

Thank you!

OpenStudy (the_fizicx99):

yw